OpenStudy (anonymous):

Which is the sixth sequence: an=1/125 (5n-1)

OpenStudy (anonymous):

a1= (which is the first term.) to find the 2nd term you plug in 2 for (an) in the formula given to you. =-4(2-1)^2 *remember an represents the term u are looking for * now you start to solve it -4(2-1)^2 -4 (1)^2 .-4^2 = 16 so your first two terms are 2 and 16 . ````````````````````````````````````````... an= -1/125 x 5^(n-1) = -1/125 x 5^( 7-1) = -1/125 x 5^6 = -1/125 x 15,625 7th = -125. *answer*

OpenStudy (anonymous):

Yea, my teacher said -125 is wrong though?

OpenStudy (anonymous):

how

OpenStudy (anonymous):